# Decomposing Multivariate Information Rates in Networks of Random Processes

### v2

@@ III.3 Formulation for Gaussian processes
the computation of (27) for each $j=1,\ldots,J$ yields the spectral MIR terms to be used in (21) for the computation of the spectral redundancy function.
Remarkably, the spectral redundancy rate defined by (21,27) is a proper redundancy function as it satisfies the axioms originally proposed by Williams and Beer, i.e. symmetry: $i^{\cap}_{Y;X_{\{\alpha_{1},\ldots,\alpha_{J}\}}}(\omega)$ is symmetric w.r.t. the $\alpha_{j},j=1,\ldots,J$; self-redundancy: $i^{\cap}_{Y;X_{\alpha}}(\omega)=i_{Y;X_{\alpha}}(\omega)$ when $|\alpha|=1$; monotonicity: $i^{\cap}_{Y;X_{\{\alpha_{1},\ldots,\alpha_{J-1},\alpha_{J}\}}}(\omega)\leq i^{\cap}_{Y;X_{\{\alpha_{1},\ldots,\alpha_{J-1}\}}}(\omega)$, with equality if $\alpha_{J-1}\subseteq\alpha_{J}$. These properties are derived straightforwardly from the minimum MIR definition (21) of the spectral redundancy rate, and from the properties of non-negativity and monotonicity of the spectral MIR (27) Nedungadi et al. (2011).
@@
We conclude discussing the practical computation of the PSD matrix (24) whose elements are exploited to estimate all the MIR terms entering the PIRD. There are several standard methods for computing the PSD matrix of multivariate time-series, including the averaged periodogram, multi taper techniques, and wavelet approaches von Sachs (2020). Here, we use the computation that induces a linear parametric representation of the observed dynamics. Specifically, the analyzed stochastic process $Z=\{Y,X_{1},\ldots,X_{M}\}$ is described as a vector autoregressive (VAR) process Lütkepohl (2005):
$Z_{n}=\sum_{k=1}^{p}\mathbf{A}_{k}Z_{n-k}+U_{n},$ (29)
@@
where $p$ is the model order, defining the maximum lag used to quantify interactions, $Z_{n}=[Y_{n}X_{1,n}\ldots X_{M,n}]^{\intercal}$ is a $M+1$-dimensional vector collecting the present state of all processes, $\mathbf{A}_{k}$ is the $(M+1)\times(M+1)$ matrix of the model coefficients relating the present with the past of the processes at lag $k$, and $U_{n}=[U_{Y,n}U_{X_{1},n}\ldots U_{X_{M},n}]^{\intercal}$ is a vector of $M+1$ zero-mean white innovations with $(M+1)\times(M+1)$ positive definite covariance matrix $\mathbf{\Sigma}_{U}=\mathbb{E}[U_{n}U_{n}^{\intercal}]$. To analyze the VAR model (29) in the frequency domain, the FT of (29) is taken to derive
$Z(\omega)=[\mathbf{I}-\sum_{k=1}^{p}\mathbf{A}_{k}e^{-\mathbf{j}\omega k}]^{-1}U(\omega)=\mathbf{H}(\omega)U(\omega),$ (30)
@@
where $Z(\omega)$ and $U(\omega)$ are the FTs of $Z_{n}$ and $U_{n}$, $\mathbf{j}=\sqrt{-1}$ and $\mathbf{I}$ is the $(M+1)$-dimensional identity matrix. The $(M+1)\times(M+1)$ matrix $\mathbf{H}(\omega)$ contains the transfer functions relating the FTs of the innovation processes in $U$ to the FTs of the processes in $Z$. Finally, the transfer matrix is exploited, together with the covariance of the VAR innovations, to derive the PSD matrix through spectral factorization Wilson (1972):
$\textbf{P}_{Z}(\omega)=\mathbf{H}(\omega)\mathbf{\Sigma}_{U}\mathbf{H}^{*}(\omega),$ (31)
@@
The practical computation of the PSD starts from the parameters $\textbf{A}_{k}$ and $\mathbf{\Sigma}_{U}$ of the VAR model (29), which can be easily estimated from realizations of the process $X$ available in the form of multivariate time series. While several procedures exist to perform VAR model identification Barnett and Seth (2014); Antonacci et al. (2020), here we use a tool Faes et al. (2012) based on classical least squares estimation Lütkepohl (2005) implemented optimizing the model order $p$ through the Akaike information criterion Akaike (1974a); remarkably, high model orders denote a non-parsimonious representation of the multivariate time series at hand, which however may indicate that the linear representation induced by the VAR model can capture at least in part possible nonlinearities in the underlying processes, provided that the fitted VAR model is stable Barnett and Seth (2015).
In closing this section we note that alternative approaches for the computation of multivariate information measures, which are not dealt with here, can rely on the model-free computation of entropy measures, which guarantees high flexibility in the representation of the dynamics but also exposes to issues related to the data-efficient estimation of entropy measures Runge et al. (2012); Vicente et al. (2011); Xiong et al. (2017); Ricci et al. (2021b). In the following sections, keeping the linear model-based approach, we will first illustrate the PIRD framework in theoretical VAR processes where the parameters are set to simulate controlled behaviors, and then apply the framework on real multivariate time series measured in the context of network physiology.
@@ IV Theoretical Examples
In this section, we characterize the behavior of the measures proposed to dissect the multivariate information shared by random processes using linear VAR models; this allows for a theoretical analysis whereby the exact profiles of the measures are computed starting from the true values imposed for the model parameters. Such parameters are varied in different simulation settings to study the time-domain and spectral behavior of the PIRD measures and to compare them with different implementations of the standard PID applied to random processes.
@@ IV.1 Effects of temporal correlations
First, we characterize the PIRD in both the time and frequency domains in simulated dynamic networks involving three processes ($Z=\{Y,X_{1},X_{2}\}$), comparing the time-domain measures with the standard PID decomposing instantaneous interactions among the processes. To do this we use a three-variate VAR process simulated with $f_{s}=1$, in which different regimes of dynamic interaction are set by varying the parameters related to zero-lag effects, lagged interactions, and autonomous dynamics. Specifically, the 3-VAR process is defined as:
$Y_{n}=cX_{1,n-1}+cX_{2,n-2}+U_{Y,n}$ (32)
$X_{1,n}=\sum_{k=1}^{2}a_{1,k}X_{1,n-k}+U_{X_{1},n}$
$X_{2,n}=\sum_{k=1}^{4}a_{2,k}X_{2,n-k}+U_{X_{2},n}$
@@
$\mathbf{\Sigma}_{U}=\begin{pmatrix}1&\sigma^{2}_{U_{YX_{1}}}&\sigma^{2}_{U_{YX_{2}}}\\ \sigma^{2}_{U_{X_{1}Y}}&1&\sigma^{2}_{U_{X_{1}X_{2}}}\\ \sigma^{2}_{U_{X_{2}Y}}&\sigma^{2}_{U_{X_{2}X_{1}}}&1\end{pmatrix},$
@@
Figure 2: Example of how the spectral redundancy rate function $i^{\cap}_{Y;X_{\{1\}\{2\}}}(f)=i^{min}_{Y;X_{1};X_{2}}(f)$ is computed as the minimum of the interaction between each individual source and the target at the specific frequency $f$, $\min\limits_{i=1,2}i_{Y;X_{i}}(f)$. The spectral PIRD allows to overcome the drawback of the MMI-PID which sets to zero one of the two unique contributions, as well as to delve into the spectral content of the investigated processes and their interactions.
@@
Figure 3: The presence of temporal correlations has a profound impact on the multivariate information shared at lag zero by multiple random processes. (a) Simulation design, where $Y$ is the target process and $\{X_{1},X_{2}\}$ is the group of sources; time-lagged interactions (solid black arrows) and zero-lag interactions (dashed gray lines) are set respectively to increase and decrease with the parameter $c\in[0-0.8]$. (b-e) Spectral profiles of the joint MIR between $Y$ and $\{X_{1},X_{2}\}$, of the individual MIR between $Y$ and $X_{1}$ and between $Y$ and $X_{2}$, and of the redundancy rate $i^{min}_{Y;X_{1};X_{2}}(f)$, obtained at varying the parameter $c$ from zero (continuous black lines) to $0.8$ (continuous pink lines). (f) Time-domain behavior of the total information shared between the target and the two sources ($I$), of the unique information shared between the target and each individual source ($\mathcal{U}_{1}$, $\mathcal{U}_{2}$), and of the redundant and synergistic information provided by the two sources to the target ($\mathcal{R}$, $\mathcal{S}$), measured using the the zero-lag PID (continuous black lines) and the PIRD (dashed gray lines) as a function of the parameter $c$.
@@
The PIRD was performed computing the spectral redundancy rate according to (21), with the spectral MIR function computed as in (27) after deriving the PSD matrix from the VAR parameters, and then obtaining the spectral PI rate through (19) which --in this case with $M=2$ sources-- yields immediately the coarse-grained terms in (20); all these terms were then integrated over the full frequency axis to obtain the unique $\mathcal{U}_{Y;X_{1}}$, $\mathcal{U}_{Y;X_{2}}$, redundant $\mathcal{R}_{Y;X_{1};X_{2}}$ and synergistic $\mathcal{S}_{Y;X_{1};X_{2}}$ information rates in the time domain. An example of how spectral redundancy is computed as the minimum MIR at each frequency $f$ is shown in Fig. 2 for one specific VAR configuration ($c$ = 0.5). Fig. 3 reports the spectral MIR and redundancy rate functions, as well as the time-domain values of the PIRD terms, investigated at varying the simulation parameter $c$. In Fig. 3f the time-domain PIRD is compared with the instantaneous PID which decomposes the MI $I(Y_{n};X_{1,n},X_{2,n})$; the latter was computed following the formulation sketched in Subsect. "Linear parametric formulation" with $T=Y_{n}$ and $S_{i}=X_{i,n}$, after deriving the zero-lag covariance of the processes via solution of the Yule Walker equations of the VAR process (the procedure is illustrated in detail in Mijatovic et al. (2024a)).
@@
The results suggest that the rate of dynamic information shared by multivariate processes is deeply affected by the balance between instantaneous and time-lagged interactions. When the three processes interact only at lag zero and do not exhibit self-dependencies ($c=0$), the spectral profiles of the MIR and redundancy rate measures are flat (Fig. 3b-e, black lines) and the time-domain PIRD and zero-lg PID measures coincide (Fig. 3f with $c=0$). Increasing the parameter $c$ determines modifications of the spectral profiles of the MIR functions (Fig. 3b-d), due to the emergence of spectral peaks around $0.1$ and $0.3$ Hz induced by the self-dependencies rising in $X_{1}$ and $X_{2}$, as well as of causal interactions along the directions $X_{1}\rightarrow Y$ and $X_{2}\rightarrow Y$. As a result, the profile of the spectral redundancy rate is also modulated by $c$ (Fig. 3e), as are the time-domain MIR and PIRD terms (Fig. 3f, dashed lines). The modification of the VAR parameters affects also the zero-lag PID (Fig. 3f, solid lines), whose information atoms were modified in a substantially different way than those of the PIRD. In fact, both redundant and synergistic contributions computed via PID ($\mathcal{R}$ and $\mathcal{S}$, solid lines in Fig. 3f) decrease towards zero at increasing $c$, while we rather expect an increase of synergy due to the emerging common child structure of the simulated system (Fig. 3a). This effect is well evidenced by the PIRD, whose synergistic contribution increases with $c$ becoming far higher than the redundant contribution ($\mathcal{R}$ and $\mathcal{S}$, dashed lines in Fig. 3f). Moreover, the PIRD detects a rise of both the unique information rates relevant to the two sources ($\mathcal{U}_{1}$ and $\mathcal{U}_{2}$, dashed lines in Fig. 3f), which is expected due to the emergence of causal interactions along the directions $X_{1}\rightarrow Y$ and $X_{2}\rightarrow Y$ with strength modulated by $c$. However, the same is not true if the unique information is measured by the zero-lag PID exploiting the time domain definition of redundancy (6) based on the MMI-PID, confirming a known limitation of such PID which always forces to zero the unique information of the source sharing the lowest information with the target (in this case $\mathcal{U}_{2}$, dashed line in Fig. 3f).

### v3

@@ III.3 Formulation for Gaussian processes
the computation of (29) for each $j=1,\ldots,J$ yields the spectral MIR terms to be used in (23) for the computation of the SMMI spectral redundancy rate.
The SMMI spectral redundancy rate defined by (23,29) is a proper redundancy function as it satisfies the axioms originally proposed by Williams and Beer Williams and Beer (2010), i.e., weak symmetry: $i^{\cap}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\{\alpha_{1},\ldots,\alpha_{J}\}}}(\omega)$ is symmetric w.r.t. the $\alpha_{j},j=1,\ldots,J$; self-redundancy: $i^{\cap}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha}}(\omega)=i_{Y;\textbf{X}_{\alpha}}(\omega)$ when $|\alpha|=1$; monotonicity: $i^{\cap}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\{\alpha_{1},\ldots,\alpha_{J-1},\alpha_{J}\}}}(\omega)\leq i^{\cap}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\{\alpha_{1},\ldots,\alpha_{J-1}\}}}(\omega)$, with equality if $\alpha_{J-1}\subseteq\alpha_{J}$; and non-negativity: $i^{\cap}_{Y;\textbf{X}_{\alpha}}(\omega)\geq 0$. These properties, which are derived from the minimum MIR definition (23) of the spectral redundancy rate, and from the properties of non-negativity and monotonicity of the spectral MIR (29) Nedungadi et al. (2011), hold equivalently for the spectral and time-domain PIRD (proofs are in Appendix A.3). Nevertheless, the SMMI measure of redundancy rate based on (29) suffers from important limitations, the main being that it is valid only for stationary Gaussian processes admitting a spectral representation Chicharro (2011), and it lacks additivity Rauh et al. (2023) in the sense that the minimum MIR of the collection of independent sets of source and target processes is not equal to the sum of the minimum MIRs evaluated separately for each set.
@@
We conclude discussing the practical computation of the PSD matrix (26) whose elements are exploited to estimate all the MIR terms entering the PIRD. There are several standard methods for computing the PSD matrix of multivariate time-series, including the averaged periodogram, multi taper techniques, and wavelet approaches von Sachs (2020). Here, we use the computation that induces a linear parametric representation of the observed dynamics. Specifically, the analyzed stochastic process ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{Z}}=\{Y,X_{1},\ldots,X_{M}\}$ is described as a vector autoregressive (VAR) process Lütkepohl (2005):
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{Z}(t_{n})}=\sum_{k=1}^{p}\mathbf{A}_{k}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{Z}(t_{n-k})}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{U}(t_{n})},$ (31)
@@
where $p$ is the model order, defining the maximum lag used to quantify interactions, $\textbf{Z}(t_{n})=[Y(t_{n})X_{1}(t_{n})\ldots X_{M}(t_{n})]^{\intercal}$ is a $M+1$-dimensional vector collecting the present state of all processes, $\mathbf{A}_{k}$ is the $(M+1)\times(M+1)$ matrix of the model coefficients relating the present with the past of the processes at lag $k$, and $\textbf{U}(t_{n})=[U_{Y}(t_{n})U_{X_{1}}(t_{n})\ldots U_{X_{M}}(t_{n})]^{\intercal}$ is a vector of $M+1$ zero-mean white innovations with $(M+1)\times(M+1)$ positive definite covariance matrix $\mathbf{\Sigma}_{\textbf{U}}=\mathbb{E}[\textbf{U}(t_{n})\textbf{U}(t_{n})^{\intercal}]$. To analyze the VAR model (31) in the frequency domain, the FT of (31) is taken to derive
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{Z}}(\omega)=[\mathbf{I}-\sum_{k=1}^{p}\mathbf{A}_{k}e^{-\mathbf{j}\omega k}]^{-1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{U}}(\omega)=\mathbf{H}(\omega){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{U}}(\omega),$ (32)
@@
where ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{Z}}(\omega)$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{U}}(\omega)$ are the FTs of $\textbf{Z}(t_{n})$ and $\textbf{U}(t_{n})$, $\mathbf{j}=\sqrt{-1}$ and $\mathbf{I}$ is the $(M+1)$-dimensional identity matrix. The $(M+1)\times(M+1)$ matrix $\mathbf{H}(\omega)$ contains the transfer functions relating the FTs of the innovation processes in U to the FTs of the processes in Z. Finally, the transfer matrix is exploited, together with the covariance of the VAR innovations, to derive the PSD matrix through spectral factorization Wilson (1972):
$\textbf{P}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{Z}}}(\omega)=\mathbf{H}(\omega)\mathbf{\Sigma}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{U}}}\mathbf{H}^{*}(\omega),$ (33)
@@
The practical computation of the PSD starts from the parameters $\textbf{A}_{k}$ and $\mathbf{\Sigma}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{U}}}$ of the VAR model (31), which can be easily estimated from realizations of the process X available in the form of multivariate time series. While several procedures exist to perform VAR model identification Barnett and Seth (2014); Antonacci et al. (2020), here we use a tool Faes et al. (2012) based on classical least squares estimation Lütkepohl (2005) implemented optimizing the model order $p$ through the Akaike information criterion Akaike (1974a); remarkably, high model orders denote a non-parsimonious representation of the multivariate time series at hand, which however may indicate that the linear representation induced by the VAR model can capture complex, potentially nonlinear dynamics of the underlying processes, provided that the fitted VAR model is stable Barnett and Seth (2015).
In closing this section we note that alternative approaches for the computation of multivariate information measures, which are not dealt with here, can rely on the model-free computation of entropy measures, which guarantees high flexibility in the representation of the dynamics but also exposes to issues related to the data-efficient estimation of entropy measures Runge et al. (2012); Vicente et al. (2011); Xiong et al. (2017); Ricci et al. (2021b). In contrast, the approach formulated in this section based on linear spectral analysis offers high computational reliability and thus favors the practical use of PIRD in a big variety of network systems encountered in real-world applications. Moreover, it is worth stressing that fitting a VAR model for computing the information decomposition measures assumes a linear model rather than a linear process, which is a subtle but important distinction: while the model has a linear structure, this does not necessarily mean that the underlying process must be linear in its dynamics. Furthermore, the presence of non-linearities in the data does not inherently imply that the fitted model will be unstable: in fact, Wold’s decomposition theorem guarantees that any stationary process admits a linear model representation, although this model may require an infinite order Hannan (1979). The critical question remains about whether the linear model applied to time series generated by a nonlinear process provides a sufficiently accurate representation of the process dynamics, and this relates to the model order that may be potentially infinite Barnett and Seth (2015, 2023). In general, choosing between the linear representation which is simpler but can be non parsimonious and a model-free approach which is theoretically well-posed but very often data demanding remains an open and interesting area for future research. In the following sections, keeping the linear model-based approach, we will first illustrate the PIRD framework in theoretical VAR processes where the parameters are set to simulate controlled behaviors, and then apply the framework on real multivariate time series measured in the context of network physiology.
@@ IV Theoretical Examples
In this section, we characterize the behavior of the measures proposed to dissect the multivariate information shared by random processes using linear VAR models; this allows for a theoretical analysis whereby the exact profiles of the measures are computed starting from the true values imposed for the model parameters. Such parameters are varied in different simulation settings to study the time-domain and spectral behavior of the PIRD measures and to compare them with different implementations of the standard PID applied to random processes. Simulations are designed to induce expected behaviors and thus aim to provide a validation of the PIRD framework in comparison to standard PID schemes.
@@ IV.1 Effects of temporal correlations
First, we characterize the PIRD in both the time and frequency domains in simulated dynamic networks involving three processes (${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{Z}}=\{Y,X_{1},X_{2}\}$), comparing the time-domain measures with the standard PID decomposing instantaneous interactions among the processes. To do this we use a three-variate VAR process simulated with $f_{s}=1$, in which different regimes of dynamic interaction are set by varying the parameters related to zero-lag effects, lagged interactions, and autonomous dynamics. Specifically, the 3-VAR process is defined as:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Y(t_{n})}=c{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X_{1}(t_{n-1})}+c{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X_{2}(t_{n-2})}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}U_{Y}(t_{n})}$ (34)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X_{1}(t_{n})}=\sum_{k=1}^{2}a_{1,k}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X_{1}(t_{n-k})}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}U_{X_{1}}(t_{n})}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X_{2}(t_{n})}=\sum_{k=1}^{4}a_{2,k}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X_{2}(t_{n-k})}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}U_{X_{2}}(t_{n})}$
@@
$\mathbf{\Sigma}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{U}}}=\begin{pmatrix}1&\sigma^{2}_{U_{YX_{1}}}&\sigma^{2}_{U_{YX_{2}}}\\ \sigma^{2}_{U_{X_{1}Y}}&1&\sigma^{2}_{U_{X_{1}X_{2}}}\\ \sigma^{2}_{U_{X_{2}Y}}&\sigma^{2}_{U_{X_{2}X_{1}}}&1\end{pmatrix},$
@@
Figure 3: Example of how the spectral redundancy rate function $i^{\cap}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\{1\}\{2\}}}(f)=i^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{SMMI}}}_{Y;X_{1};X_{2}}(f)$ (dashed pink line) is computed as the minimum of the interaction between each individual source and the target at the specific frequency $f$, $\min\limits_{i=1,2}i_{Y;X_{i}}(f)$, where the spectral profiles $i_{Y;X_{1}}(f),i_{Y;X_{2}}(f)$ are arbitrarily depicted as solid black and grey lines, respectively. The spectral PIRD allows to overcome the drawback of the MMI-PID which sets to zero one of the two unique contributions, as well as to delve into the spectral content of the investigated processes and their interactions.
@@
Figure 4: The presence of temporal correlations has a profound impact on the multivariate information shared at lag zero by multiple random processes. (a) Simulation design, where $Y$ is the target process and $\{X_{1},X_{2}\}$ is the group of sources; time-lagged interactions (solid black arrows) and zero-lag interactions (dashed grey lines) are set respectively to increase and decrease with the parameter $c\in[0-0.8]$. (b-e) Spectral profiles of the joint MIR between $Y$ and $\{X_{1},X_{2}\}$, of the individual MIR between $Y$ and $X_{1}$ and between $Y$ and $X_{2}$, and of the redundancy rate $i^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{SMMI}}}_{Y;X_{1};X_{2}}(f)$, obtained at varying the parameter $c$ from zero (continuous black lines) to $0.8$ (continuous pink lines). (f) Time-domain behavior of the total information shared between the target and the two sources ($I$), of the unique information shared between the target and each individual source ($\mathcal{U}_{1}$, $\mathcal{U}_{2}$), and of the redundant and synergistic information provided by the two sources to the target ($\mathcal{R}$, $\mathcal{S}$), measured using the the zero-lag PID (continuous black lines) and the PIRD (dashed grey lines) as a function of the parameter $c$.
@@
The PIRD was performed computing the SMMI redundancy rate according to (23), with the spectral MIR function computed as in (29) after deriving the PSD matrix from the VAR parameters, and then obtaining the spectral PI rate through (21) which --in this case with $M=2$ sources-- yields immediately the coarse-grained terms in (22); all these terms were then integrated over the full frequency axis to obtain the unique $\mathcal{U}_{Y;X_{1}}$, $\mathcal{U}_{Y;X_{2}}$, redundant $\mathcal{R}_{Y;X_{1};X_{2}}$ and synergistic $\mathcal{S}_{Y;X_{1};X_{2}}$ information rates in the time domain. An example of how spectral redundancy is computed as the minimum MIR at each frequency $f$ is shown in Fig. 3 for one specific VAR configuration ($c$ = 0.5). Fig. 4 reports the spectral MIR and redundancy rate functions, as well as the time-domain values of the PIRD terms, investigated at varying the simulation parameter $c$. In Fig. 4f the time-domain PIRD is compared with the instantaneous PID which decomposes the MI $I(Y(t_{n});X_{1}(t_{n}),X_{2}(t_{n}))$; the latter was computed following the formulation sketched in Subsect. "Linear parametric formulation" with $T=Y(t_{n})$ and $S_{i}=X_{i}(t_{n})$, after deriving the zero-lag covariance of the processes via solution of the Yule Walker equations of the VAR process (the procedure is illustrated in detail in Mijatovic et al. (2024a)).
@@
The results suggest that the rate of dynamic information shared by multivariate processes is deeply affected by the balance between instantaneous and time-lagged interactions. When the three processes interact only at lag zero and do not exhibit self-dependencies ($c=0$), the spectral profiles of the MIR and redundancy rate measures are flat (Fig. 4b-e, black lines) and the time-domain PIRD and zero-lag PID measures coincide (Fig. 4f with $c=0$). Increasing the parameter $c$ determines modifications of the spectral profiles of the MIR functions (Fig. 4b-d), due to the emergence of spectral peaks around $0.1$ and $0.3$ Hz induced by the self-dependencies rising in $X_{1}$ and $X_{2}$, as well as of causal interactions along the directions $X_{1}\rightarrow Y$ and $X_{2}\rightarrow Y$. As a result, the profile of the spectral redundancy rate is also modulated by $c$ (Fig. 4e), as are the time-domain MIR and PIRD terms (Fig. 4f, dashed lines). The modification of the VAR parameters affects also the zero-lag PID (Fig. 4f, solid lines), whose information atoms were modified in a substantially different way than those of the PIRD. In fact, both redundant and synergistic contributions computed via PID ($\mathcal{R}$ and $\mathcal{S}$, solid lines in Fig. 4f) decrease towards zero at increasing $c$, while we rather expect an increase of synergy due to the emerging common child structure of the simulated system (Fig. 4a). This effect is well evidenced by the PIRD, whose synergistic contribution increases with $c$ becoming far higher than the redundant contribution ($\mathcal{R}$ and $\mathcal{S}$, dashed lines in Fig. 4f). Moreover, the PIRD detects a rise of both the unique information rates relevant to the two sources ($\mathcal{U}_{1}$ and $\mathcal{U}_{2}$, dashed lines in Fig. 4f), which is expected due to the emergence of causal interactions along the directions $X_{1}\rightarrow Y$ and $X_{2}\rightarrow Y$ with strength modulated by $c$. However, the same is not true if the unique information is measured by the zero-lag PID exploiting the time domain definition of redundancy (17) based on the MMI-PID, confirming a known limitation of such PID which always forces to zero the unique information of the source sharing the lowest information with the target (in this case $\mathcal{U}_{2}$, dashed line in Fig. 4f).